OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -2/5x and contains the point (5,-3) 2)Write the equation of a line that is perpendicular to the line y = 1/4x + 2 and contains the point (11,4).

OpenStudy (anonymous):

parallel means same slope slope of given line is -2/5 so we know the new slope is also -2/5 plug in the provided points:\[(-3)={-{2\over5}}(5)+b\]\[-3=2+b\]\[-1=b\]now write the equation with the following information:\[m=-{2\over5}\]\[b=-1\]\[y=mx+b\]

OpenStudy (anonymous):

y=-2/5-1?

OpenStudy (anonymous):

@yummydum

OpenStudy (anonymous):

dont forget your x :)\[y=-{2\over5}x-1\]

OpenStudy (anonymous):

omg thanks tats the first real help ive gotten ! ur helping me learn !! @yummydum

OpenStudy (anonymous):

YAY ^_^

OpenStudy (anonymous):

glad you're learning, makes me happy :)

OpenStudy (anonymous):

can u explain number 2 and i solve it ? @yummydum

OpenStudy (anonymous):

perpendicular lines have opposite, reciprocal slopes lets say a line has slope 3/2, its perpendicular line would have slope -2/3 and after you find what the perpendicular slope is, follow the same steps, plug in the points, write the equation and tada! :)

OpenStudy (anonymous):

so if your slope is 1/4, what would be the perpendicular slope?

OpenStudy (anonymous):

y=-4/1x-1

OpenStudy (anonymous):

\[(4)=-4(11)+b\]\[4=-44+b\]\[48=b\]\[y=-4x+48\]

OpenStudy (anonymous):

remember that the x you plug in is 11 and the y you plug in is 4 the slope you use is -4 and you have to find b so after solving 4=-4(4)+b you get b=48 thus your equation would be y=-4x+48 :)

OpenStudy (anonymous):

i meant after solving 4=-4(11)+b you get b=48

OpenStudy (anonymous):

thanks it makes more since now i forgot to plug in

OpenStudy (anonymous):

lol i thought so :)

OpenStudy (anonymous):

hey do u think u can help with my algebraic proofs question i asked a couple days ago ? @yummydum

OpenStudy (anonymous):

umm....is it okay if i help you tomorrow? :D its 3am my time and i have to study for an exam that starts in just about 4-5 hrs.. yeah

OpenStudy (anonymous):

ok !!!! lol